### v1

# AVTENet: Audio-Visual Transformer-based Ensemble Network Exploiting Multiple Experts for Video Deepfake Detection

Abstract
Forged content shared widely on social media platforms is a major social problem that requires increased regulation and poses new challenges to the research community. The recent proliferation of hyper-realistic deepfake videos has drawn attention to the threat of audio and visual forgeries. Most previous work on detecting AI-generated fake videos only utilizes visual modality or audio modality. While there are some methods in the literature that exploit audio and visual modalities to detect forged videos, they have not been comprehensively evaluated on multi-modal datasets of deepfake videos involving acoustic and visual manipulations. Moreover, these existing methods are mostly based on CNN and suffer from low detection accuracy. Inspired by the recent success of Transformer in various fields, to address the challenges posed by deepfake technology, in this paper, we propose an Audio-Visual Transformer-based Ensemble Network (AVTENet) framework that considers both acoustic manipulation and visual manipulation to achieve effective video forgery detection. Specifically, the proposed model integrates several purely transformer-based variants that capture video, audio, and audio-visual salient cues to reach a consensus in prediction. For evaluation, we use the recently released benchmark multi-modal audio-video FakeAVCeleb dataset. For a detailed analysis, we evaluate AVTENet, its variants, and several existing methods on multiple test sets of the FakeAVCeleb dataset. Experimental results show that our best model outperforms all existing methods and achieves state-of-the-art performance on Testset-I and Testset-II of the FakeAVCeleb dataset.
Index Terms:
Deepfake detection, Video forgery detection, Audio-visual feature fusion, Transformer, AVTENet
I Introduction
Figure 1: An example of audio-visual deepfakes. Cyber-attackers aim to create convincing counterfeit videos by manipulating audio and visual streams. Sharing such malicious content can have harmful consequences. Therefore, our proposed method exploits manipulations in both audio and visual streams to effectively detect forged videos.
Advances in social media and machine learning technology have made it easier to generate and spread quickly hyper-realistic counterfeit content, commonly known as Deepfakes. Deepfake refers to forged media content in which the face and/or voice of the source person has been manipulated with the face and/or voice of a target person using deep learning algorithms. The rapid development of modern deep learning technology is driving multimedia processing, achieving incredible results in transforming multimedia content (images, audio, and video). Nowadays, it is a cinch to create or manipulate widely available online media content through various advanced tools and techniques; on top of that, the availability of large-scale datasets grants supplementary potential to deepfake generative models. Models such as Generative Adversarial Networks (GANs) [1] and Variational Autoencoders (VAEs) [2] can create highly realistic synthetic media content that is indistinguishable. Face Swapping, Face Reenactment, Face Editing, and Face Synthesis are different categories of deepfake techniques that help produce convincing and well-crafted facial deepfake media [3], with face swapping being the most common visually manipulated deepfake today. Similarly, Text-to-Speech Synthesis (TTS) and Voice Conversion (VC) are audio deepfake generation techniques that enable deception by cloning an individual’s voice. In addition to deepfake audio generation, automatic speaker verification (SV) systems help recognize the speaker’s identity or identify the speaker in the audio [4, 5]. It has been observed that existing SV systems are easily spoofed by morphing the audio signal. State-of-the-art (SOTA) deepfake generative models have a wide range of beneficial applications [6]; however, deepfakes can have potentially serious adverse effects and can pose a threat when used for malicious purposes, such as revenge porn, cyber-crime, misinformation, political denigration, privacy, legal and ethical issues. Misinformation or disinformation [7] spread through deepfakes can lead to chaos or political instability [8], while cyber-attacks can escalate the threat of access to sensitive information. Prevalent misinformation on online social media leads to the spread of unverified rumors and can lead to social disputes. Therefore, effective deepfake detection methods are urgently needed to mitigate these potential harms, and video deepfake detection has become a focus for media forensics researchers.
At first instance, deepfake content involves uni-modal manipulation in audio or vision. But more recently, multi-modal manipulation has been noticed in generating more realistic deepfake videos. The viral Instagram video of Mark Zuckerberg (Meta’s CEO), as shown in Fig. 1, is the ultimate example of audio-visual deepfakes that combine acoustic and visual forgeries. Detecting such convincing deepfake videos is a crucial but challenging task. To combat the trend of deepfakes, several deep learning-based solutions have been proposed to detect forgeries in videos. However, most existing models only focus on uni-modal detection of video deepfakes, and lack evaluation on multi-modal deepfake datasets. Uni-modal methods limit the performance of detectors to certain scenarios; for example, a video-only detector cannot detect acoustic manipulation, while a detector that relies solely on the audio modality can easily be fooled by visual manipulation. Moreover, the lack of benchmark multi-modal datasets that contain both acoustic and visual forgeries hinders the training of multi-modal video deepfake detectors, as existing video datasets overlook audio deepfakes and multi-modal deepfakes. Therefore, to address the above issues, our study focuses on an audio-visual approach evaluated on a multi-modal deepfake dataset to take a step toward applications in real-world scenarios.
Figure 2: Illustrations of the audio and visual streams of some samples selected from the FakeAVCeleb Dataset, where real streams are indicated in green, and fake ones are indicated in red.
Recently, transformers [9] have emerged as a promising alternative and marked a dominant shift from Convolutional Neural Networks (CNNs) in computer vision tasks due to their ability to model long-range dependencies and capture global context. The sequential processing nature of CNN limits the network’s ability to capture long-range dependencies; however, the attention mechanism in the transformer allows it to model long-range global context and relationships. Existing methods mainly use CNN-based solutions to deal with deepfake detection [10, 11, 12, 13, 14, 15, 16, 17]. These CNN-based methods process input data sequentially, while transformers can process it efficiently by processing the entire input sequence in parallel. Inspired by transformers’ success in various fields, we intend to apply it to the video forgery detection task by leveraging acoustic and visual features. We believe that multiple modalities provide complementary information and stable inferences. Therefore, to fill the above gap, we propose AVTENet, a novel Audio-Visual Transformer-based Ensemble Network that utilizes both acoustic and visual features for effective video forgery detection. The main contributions of our work are summarized as follows:
  • We propose an audio-visual transformer-based ensemble network, AVTENet, demonstrating that audio-only, video-only, and audio-visual transformer-based networks can be effectively combined to improve robustness and detection accuracy. The proposed approach can address scenarios where only a single modality or both modalities in audio vision are manipulated.
  • To understand the strengths and weaknesses of various ensemble approaches, we present four variants of AVTENet and show that the feature fusion approach performs better than majority voting, score fusion, and average score fusion.
  • We experimentally confirm the effectiveness of AVTENet and empirically demonstrate that our transformer-based ensemble model outperforms the previous CNN-based ensemble network [10].
  • To evaluate the proposed method, we use the FakeAVCeleb dataset [18], which is a multi-modal audio-visual deepfake dataset. We conduct a comprehensive experimental comparison between the proposed method and existing single-modal and multi-modal methods. The results confirm that our proposed method achieves the highest performance on the test sets of the FakeAVCeleb dataset, outperforming all compared methods.
The remainder of this paper is organized as follows. Sec. II reviews related work. Sec. III illustrates the proposed methodology. Sec. IV outlines the experimental setup and reports the results. Finally, Sec. V provides conclusions and future work.
II Related Work
II-A Video Deepfake Generation
Deepfake technology has become more sophisticated as it uses artificial intelligence (AI) techniques to generate, manipulate, or alter video content, thereby adding significant threats to society, legislation, individuals, community consensus, journalism, and cyber-security. Face2Face [19], Neural Talking-Heads [20], First Order Motion Model for Image Animation [21], and Deep Video Portraits [22] are some of the deepfake generation techniques. Likewise, there exist different types of deepfake videos, including lip-syncing, face-swapping, and voice manipulation. Generative Adversarial Networks (GANs) [23] become a popular deep-learning method to generate realistic images or videos. For instance, in [20] and [24] GAN-based video deepfake generation methods were proposed. The method proposed in [24] generates a dance video of a person from a single image. First, a sequence of human poses is generated and rendered in the video. To generate realistic animated dance videos, this method combines pose estimation, key-point detection, and image synthesis. Similarly, the few-shot learning method proposed in [20] uses a few reference images to generate realistic videos of talking heads. Their method creates a realistic video sequence of a target person speaking by mapping a number of reference images. Another GAN-based method for image-to-image translation was proposed in [25]. To ensure high perceptual qualities of the generated images, the model was trained with a perceptual loss function. The authors evaluated their method on several image generation tasks, such as image colorization, painting, and super-resolution.
II-B Video Deepfake Detection
II-B1 Uni-modal Deepfake Detection
Previous work has mainly focused on a single modality to detect forgeries in videos since synthetic videos generally have less manipulation in the audio track. Some researchers have worked to detect forgeries in videos by exploiting visual artifacts such as inconsistent facial expressions and movements [26, 27], irregular head movements or angles [28, 26], mismatched shadows or lightening [29], distorted or blurred edges around the face or body [30], unnatural eye blinking or movement [31], unnatural background or scenery [32]. The authors of [33] argue that several facial editing techniques exhibit visual artifacts in videos that can be tracked and used to detect video forgeries. Considering this rationale, they demonstrated that even simple visual cues can reveal deepfakes. The authors of [31] used eye blinking, a physiological signal that is not well presented in synthetic videos, to detect fake face videos. Other researchers have exploited emotional inconsistency [26], synchronization between speech and lip movements [11], or motion analysis [34]. A biometrics-based forensics technique was proposed in [35] to detect faceswap deepfakes by exploiting appearance and behavior, especially facial expressions and head movements. MesoNet and MesoInception are two popular models proposed in [12]. Both are compact CNN-based models with a few layers that exploit mesoscopic properties to detect facial tampering in videos. MesoNet analyzes the residual high-frequency content of video frames. XceptionNet [13] is built on depth-wise separable convolutional layers. The authors of [14] and [15] also proposed other CNN-based methods for video forgery detection. The method proposed in [14] employs composite coefficients that effectively scale all dimensions of depth, width, and resolution uniformly. The method proposed in [15] utilizes audio and video streams to extract features and then combines these features to better detect deepfake videos. A transformer network based on an incremental learning strategy is proposed in [36], which utilizes face images and their UV texture generated by a 3D face construction method using a single-face input image to detect synthetic videos. The same authors later proposed a hybrid transformer network to detect forged videos [37]. The transformer network is trained end-to-end using XceptionNet and EfficientNet-B4 as feature extractors.
II-B2 Multi-modal Deepfake Detection
Considering that the auditory context plays a crucial role in visual object perception [38], few recent attempts have addressed the challenge of detecting multi-modal deepfake videos using multiple modalities such as video, audio, and text. For instance, inspired by the Siamese networks and triplet loss, a learning-based approach is proposed in [39], which uses acoustic and visual features of the same video to enhance the information for learning. The authors extracted affective cues and used them to compare the emotions perceived in the two modes to discern whether the video was genuine or manipulated. The authors of [40], [41] and [42] combined audio and visual modalities to develop a video deepfake detector that can handle audio-only, video-only, and audio-visual manipulations. In [40], a promising approach is proposed to exploit the intrinsic synchronization between audio and visual modalities to potentially detect increasingly prevalent deepfakes. In [41], an audio-visual person-of-interest deepfake detector that exploits specific biometrics of people is proposed to cope with the variety of manipulation methods used in synthetic media generation. The high generalizability of the detector is due to its independence from any manipulation techniques and training on videos of real talking faces. Similarly, for audiovisual forgery detection, authors in [42] incorporate audio loss, video loss and audiovisual loss to capture the inconsistency of multi-modality and artifacts from individual modalities. Additionally, ensemble approaches were used in [43] and [10] to improve the accuracy of deep fake detection. In [43], incorporating supplementary textual features has been shown to be beneficial in effectively handling multi-modal fake content detection. The authors built an ensemble network combining uni-modal and cross-modal classifiers to distinguish forged from genuine video clips. On the other hand, the ensemble method in [10] combines CNN-based audio-only, video-only, and audio-visual networks to detect multi-modal deepfakes. Furthermore, the authors of [16] argue that the manipulation of audio or visual flow in video can lead to a lack of harmony between them, such as speech-lip non-synchrony and unnatural facial movements. Therefore, they utilized modality dissonance scores to measure the dissimilarity between two modalities to identify genuine or spoofed videos. Another deep learning method that exploits the audio and visual modalities and considers the importance of speech-lip synchronization for the task of video forgery detection is proposed in [11]. Their audio-visual deepfake detector checks the synchrony between the lip sequence extracted from the video and the synthetic lip sequence generated from the audio track. Similarly, authors in [44] adopted a self-supervised transformer-based approach that utilizes contrastive learning. Their method allows paired acoustic and visual streams to learn mouth motion representations propelling paired representation to be closer and unpaired to stay farther in order to determine audio-visual forgery. Another self-supervised learning approach in [45] exploits an audio-visual temporal synchronization by evaluating consistency between the acoustic stream and faces in a video clip to determine forgery. Although uni-modal methods have been extensively studied in video forgery detection, the underlying correlations implicit between different perception modes have not been fully explored and exploited.
II-C Multi-modal Deepfake Datasets
To the best of our knowledge, DFDC [46] and FakeAVCeleb [18] are the only multi-modal datasets that include both acoustic and visual manipulations in videos.
The DFDC dataset [46] is larger than the FakeAVCeleb dataset, containing 100,000 video clips collected from 3,426 paid subjects. DFDC provides a benchmark for evaluating the performance of deepfake detection models and serves as a crucial resource for advancing research in the field of deepfake detection. It has been released as part of a competition to promote the development of deepfake detection methods. The deepfake videos in this dataset were generated using various deepfake techniques, such as GAN-based methods and non-learning methods. It consists of real and fake videos and considers both acoustic and visual manipulations.
The FakeAVCeleb dataset [18] is a multi-modal audio-visual dataset released after DFDC. The dataset was carefully generated through multiple manipulation methods, taking into account the balance among ethnic backgrounds, gender, and age groups. The real videos of 500 celebrities originating from the Voxceleb2 dataset [47] collected from YouTube form the base set. This base set was further used to generate 20,000 deepfake videos through manipulation techniques, including Faceswap [48], FSGAN [49], Wav2Lip [50], and RTVC (real-time-voice-cloning) [51].
Based on its advantages mentioned above, we use the FakeAVCeleb dataset to evaluate our multi-modal deepfake detector AVTENet. Fig. 2 shows a few samples of each category from the FakeAVCeleb dataset. We do not use the DFDC dataset due to its extreme environmental settings, e.g., some instances are poorly illuminated or overexposed, or sometimes the subject’s face may not be facing the camera. Another reason is that its labels lack separate labels for the acoustic and visual modes in the video.
Figure 3: Overview of the proposed AVTENet model for detecting video forgeries by simultaneously utilizing acoustic and visual cues. It consists of a video-only network (VN), an audio-only network (AN), and an audio-visual network (AVN), each of which is an independent transformer-based classifier, and a decision-making module (DM) that produces final predictions based on the three classifiers through various ensemble strategies.
III Proposed Method
The natural auditory context provides distinctive, independent, and diagnostic information about the visual world, which directly impacts the perceptual experience of visual objects [38]. To combat the emerging threat of video deepfakes, we leverage both acoustic and visual information to detect forgeries in videos. We use the transformer-based model because its self-attention mechanism can catch inconsistencies or manipulations to identify video forgeries. Furthermore, we employ ensemble learning to exploit the fusion of complementary information, modality-specific patterns, and multi-modal features to improve the robustness of the detection system. Accordingly, we propose an audio-visual transformer-based ensemble network (AVTENet) that fully exploits acoustic and visual information to detect video forgeries by leveraging modality-specific cues and complementary insights of joint audio and visual information. Fig. 3 shows an overview of our proposed AVTENet model, which consists of three key networks, namely a video-only network (VN), an audio-only network (AN), and an audio-visual network (AVN), as well as a decision-making module (DM). The three key networks are different-modality transformer-based networks integrated with pre-trained models through self-supervised learning (SSL) or supervised learning (SL). Then, DM integrates the outputs of these three classifiers according to different fusion strategies. Given a test video $x$, the decision is made according to
${{\mathrm{AVTENet}(x)}={\mathrm{DM}(C_{v}(x)},{C_{a}(x)},{C_{av}(x))}},$ (1)
where $C_{v}$, $C_{a}$, and $C_{av}$ denote the video-only, audio-only, and audio-visual classifiers, respectively, and $\mathrm{DM}$ denotes the function of decision making.
III-A Video-only Network (VN)
The VN module extracts relevant spatiotemporal features from video frames through its self-attention mechanism to capture spatiotemporal patterns and long-range dependencies, and then outputs a vector representation that captures essential visual information.
Inspired by the remarkable progress achieved by the video vision transformer ViViT [52] in video classification tasks, we seamlessly integrate it into the visual backbone of AVTENet. As shown in the top part of Fig. 3, we use a factorized encoder model that consists of two separate transformer encoders. The first is a spatial encoder, which delves into the interaction among tokens extracted from the same temporal index. The second is a temporal encoder that models interactions among tokens from different temporal indices. We first divide the input video into small segments, a series of patches is extracted from each segment in the form of a tubelet and passed through the spatial encoder to produce an encoding vector representing the segment. The output encoding vectors serve as input to the temporal transformer along with an additional classification (cls) token to extract the final representation of the entire video.
For VN training, a dataset $D^{v}=\{v_{i},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $v_{i}$ denotes the video stream of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $v_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the ViViT model extracts the single-vector representation $E_{v}$ (the $cls$ token of the temporal transformer) of each training video clip. A linear layer is used as the classifier. The ViViT model is pre-trained on the Kinetics dataset [52]. During the VN training process, not only is the linear classification layer trained, but also the pre-trained ViViT model is fine-tuned. The model is trained with the binary cross-entropy loss defined as
${\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{n}\sum_{i=1}^{n}y_{i}\log(\hat{y}_{i})+\left(1-y_{i}\right)\log\left(1-\hat{y}_{i}\right)},$ (2)
where $n$ is the total number of training samples, and $\hat{y}_{i}$ denotes the predicted score.
In the inference phase of VN, given a test sample $x$, the prediction is conducted by
$C_{v}(x)=\mathrm{VN}(x_{v}),$ (3)
where $x_{v}$ is the video stream (i.e., the image sequence) of $x$. Note that in the inference phase of our ensemble model, the DM module utilizes either the single-vector representation or the predicted score corresponding to the “fake” of the video-only network.
III-B Audio-only Network (AN)
The AN module takes as input an audio spectrogram, where time and frequency elements help learn and capture acoustic patterns, temporal dynamics, and other audio-specific characteristics, and returns a high-dimensional feature representation.
Transformers perform incredibly well in audio processing as their self-attention mechanisms allow for capturing long-range dependencies in audio. In this work, we use the Audio Spectrogram Transformer (AST) [53] as the audio backbone of AVTENet. As shown at the bottom of Fig. 3, the AN module takes as input a spectrogram that is further divided into a sequence of 16x16 patches with an overlap of 6 in both the time and frequency dimensions. A linear projection layer transforms each patch into a 1D embedding. Since the patch sequence lacks input order information, trainable positional embeddings are incorporated into each patch embedding to enable capturing spatiotemporal structure from a 2D audio spectrogram. Additionally, a classification (cls) token is added to the sequence. The output cls embedding of the transformer encoder encapsulates information about characteristics that can be used to determine challenging and deceptive alternations in the audio.
Figure 4: Four variants of the proposed AVTENet model for video forgery detection based on different ensemble strategies: (a) $AVTENet_{mv}$ based on majority voting, (b) $AVTENet_{asf}$ based on average score fusion, (c) $AVTENet_{sf}$ based on score fusion, and (d) $AVTENet_{ff}$ based on feature fusion.
For training AN, a dataset $D^{a}=\{a_{i},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $a_{i}$ denotes the audio track of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $a_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the AST model is used to extract the single vector representation $E_{a}$ (the $cls$ token) of the mel-spectrogram of each training audio clip. A linear layer is used as the classifier. The AST model is pre-trained on Audioset [53]. During the training process of AN, the linear classification layer is trained, and the pre-trained AST model is fine-tuned. As with the video-only network, the cross-entropy loss is used to train the audio-only network.
In the inference phase of AN, given a test sample $x$, the prediction is conducted by
$C_{a}(x)=\mathrm{AN}(x_{a}),$ (4)
where $x_{a}$ is the audio track of $x$. As with the video-only network, in the inference phase of our ensemble model, either the single-vector representation or the predicted score corresponding to “fake” of the audio-only network is used by the DM module.
III-C Audio-Visual Network (AVN)
The AVN module exploits both visual and acoustic information by jointly learning paired audio and visual inputs that provide complementary information and capture meaningful representations between these distinct modalities to expose fabrication in videos.
We use the self-supervised learning framework AV-HuBERT [54] as the audio-visual backbone of AVTENet. As shown in the middle of Fig. 3, the AVN module simultaneously uses the audio spectrogram and the video frame stack of the lip region extracted from the video as inputs to the lightweight modality-specific encoders. The frames of the lip region are passed through a ResNet-based feature extractor to extract relevant visual features. At the same time, the audio spectrogram is passed through the feed-forward network (FFN) to extract the acoustic features. The visual and acoustic features are then fused and fed to a shared transformer encoder to extract jointly contextualized audio-visual representations that encapsulate the correlation between acoustic and visual modalities. These audio-visual representations are further passed through a Multiscale Temporal Convolutional Network (MS-TCN) and a temporal pooling layer to produce a single-vector representation.
For training AVN, a dataset $D^{av}=\{a_{i},v_{i}^{lip},y_{i}\}_{i=1}^{n}$ is extracted from the training set of the FakeAVCeleb dataset, where $a_{i}$ and $v_{i}^{lip}$, respectively, denote the audio track and the lip image sequence of the $i$-th training sample $x_{i}$, and $y_{i}$ is the label for $x_{i}$ (0 for fake and 1 for real). As shown in Fig. 3, the AV-HuBERT model, MSTCN network, and temporal pooling layer are used to extract the vector representation $E_{av}$ of the mel-spectrogram of the audio track and the lip-image sequence of each training video clip. A linear layer is used as the classifier. The AV-HuBERT model is pre-trained on the LRS3 dataset [55]. During AVN’s training process, not only MSTCN with linear classification layer is trained, but also the pre-trained AV-HuBERT model is fine-tuned. As with the video-only and audio-only networks, the cross-entropy loss is used to train the audio-visual network.
In the inference phase of AVN, given a test sample $x$, the prediction is conducted by
$C_{av}(x)=\mathrm{AVN}(x_{a},x_{v^{lip}}),$ (5)
where $x_{a}$ and $x_{v^{lip}}$ are the audio track and the lip-image sequence of $x$, respectively. As with the video-only and audio-only networks, in the inference phase of our ensemble model, either the single-vector representation or the predicted score corresponding to “fake” of the audio-visual network is used by the DM module.
In addition to the above AV-HuBERT-based AVN setup, we also study another setup. The audio and video embeddings are extracted via AST (same as AN) and ViViT (same as VN). The corresponding $cls$ tokens are concatenated as $E_{av}$. The prediction is also conducted by Eq. (5), but using ($x_{a}$, $x_{v}$) instead of ($x_{a}$, $x_{v^{lip}}$) as input. An overview of the AVTENet model based on this AVN setting is shown in Fig. 5. Compared with the AVTENet model in Fig. 3, the only difference is that the AV-HuBERT-based AVN is replaced by the AST&ViViT-based AVN.
Figure 5: Overview of another AVTENet model that consists of AST-based AN, ViViT-based VN, and AST&ViViT-based AVN.
III-D Decision-Making Module (DM):
As shown in Fig. 3, the DM module of AVTENet is used to integrate the outputs of VN, AN, and AVN to produce final predictions that indicate potential forgery in the video. In this work, we investigate several fusion strategies.
III-D1 Majority Voting
As shown in Fig. 4(a) and Algorithm 1, the DM module outputs “fake” if at least two component classifiers (VN, AN, and AVN) output “fake”. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{mv}$. Note that no additional models need to be trained for majority voting fusion.
Algorithm 1 Majority Voting Algorithm for $\mathrm{AVTENet}_{mv}$
0:  $P_{v},P_{a}$, and $P_{av}\left(P_{i}\in 0\right.$ or 1))))
0:  AVTENet_Prediction $P\left(P\in 0\right.$ or 1))))
1:  Initialize $vote\_count$ variable, $vote\_count$ = $P_{v}+P_{a}+P_{av}$
2:  if  $vote\_count$ $\geq 2$ then
3:     $P$ = 1
4:  else
5:     $P$ = 0
6:  end if
7:  return  $P$
III-D2 Average Score Fusion
As shown in Fig. 4(b), the DM module outputs “fake” if the average of the output scores of the three component classifiers (VN, AN, and AVN) exceeds a preset threshold. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{asf}$. Note that no additional models need to be trained for average score fusion.
III-D3 Score Fusion
As shown in Fig. 4(c), in score fusion, the output scores of the three component classifiers (VN, AN, and AVN) are concatenated as input to a linear layer trained on the same training data as the three component classifiers. When training the linear layer, the parameters of the three component classifiers are fixed. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{sf}$.
III-D4 Feature Fusion
As shown in Fig. 4(d), in feature fusion, the output representations of the penultimate layers of the three component classifiers (VN, AN, and AVN) are concatenated as input to a linear layer trained on the same training data as the three component classifiers. When training the linear layer, the parameters of the three component classifiers are fixed. The corresponding AVTENet model is termed $\mathrm{AVTENet}_{ff}$.
IV Experimental Results
This section presents the experiment setup, including the dataset, data analysis and preprocessing, evaluation metrics, training hyperparameters, and experimental results.
IV-A Dataset
Our experiments are conducted on the FakeAVCeleb dataset [18], an audio-visual dataset released in 2022 specifically designed for the deepfake detection task. It is based on a collection of 500 YouTube videos featuring 500 ethnically diverse celebrities from various fields, including music, film, sports, and politics, taking into account age, geography, race, and gender diversity. Four latest deepfake manipulation and synthetic speech generation techniques, including Faceswap [48], FSGAN [49], Wav2Lip [50], and Real-Time-Voice-Cloning (RTVC) [51], and their combinations were used to generate 19,500 fake video samples from these 500 videos. As a result, the dataset contains a total of 20,000 video samples. The videos corresponding to 70 celebrities are used for testing, and the videos corresponding to the remaining 430 celebrities are used for training. These video samples are divided into four categories according to whether the audio and video modalities are manipulated, namely RealVideo-RealAudio ($R_{v}R_{a}$), RealVideo-FakeAudio ($R_{v}F_{a}$), FakeVideo-RealAudio ($F_{v}R_{a}$) and FakeVideo-FakeAudio ($F_{v}F_{a}$), where each video is carefully labeled to indicate its authenticity and facilitate training and testing forgery detection models.
TABLE I: Manipulation modalities in each test set. $\checkmark$ indicates manipulation, while $X$ indicates no manipulation.
Testsets Audio Manipulation Video Manipulation
Testset-1 $\checkmark$ $\checkmark$
Testset-II $\checkmark$ $\checkmark$
faceswap $X$checkmark
faceswap-wav2lip $\checkmark$ $\checkmark$
fsgan $X$checkmark
fsgan-wav2lip $\checkmark$ $\checkmark$
RTVC $\checkmark$ $X$
wav2lip $\checkmark$ $\checkmark$
IV-A1 Testing Sets
To conduct a comprehensive evaluation of various uni-modal (AN and VN) and multi-modal (AVN and AVTENet) detection methods using the FakeAVCeleb dataset, we used eight different test sets, including Testset-I, Testset-II, faceswap, faceswap-wav2lip, fsgan, fsgan-wav2lip, RTVC, and wav2lip. Except for Testset-I and Testset-II, each test set contains the same set of 70 genuine videos of 70 subjects unseen in training and a different set of 70 fake videos generated using specific manipulation techniques. Testset-I and Testset-II are the main evaluation test sets, which also contain the same 70 genuine videos of 70 subjects unseen in training and 70 fake videos. However, Testset-I contains the same number of fake samples from each manipulation technique, while Testset-II contains an equal number of fake samples from the $R_{v}F_{a}$, $F_{v}F_{a}$ and $F_{v}R_{a}$ categories. Table I summarizes the manipulation modalities involved in each test set. Multiple test sets can provide an exhaustive understanding of the strengths and limitations of AVTENet and individual classifiers on fake video samples generated using different deepfake techniques.
IV-A2 Training Sets for Different Models
In the experiments, the training data are video samples corresponding to 430 subjects in FakeAVCeleb. However, training uni-modal and multi-modal classifiers requires specific data settings, so we customize the dataset according to the requirements of each network. Table II shows the training data setting for each network according to the categorical labels of FakeAVCeleb.
TABLE II: Training data settings for different networks.
Classifier Class Category
Video Network Fake $F_{v}F_{a}$
$F_{v}R_{a}$
Real $R_{v}F_{a}$
$R_{v}R_{a}$
Audio Network Fake $R_{v}F_{a}$
$F_{v}F_{a}$
Real $F_{v}R_{a}$
$R_{v}R_{a}$
Audio-Visual Network Fake $F_{v}F_{a}$
$F_{v}R_{a}$
$R_{v}F_{a}$
Real $R_{v}R_{a}$
Video-only Network (VN): The VN classifier is trained to exploit manipulations in visual content. When constructing the training set for VN, as shown in Table II, the fake class includes training samples in the $F_{v}F_{a}$ and $F_{v}R_{a}$ categories because the visual content in these categories is manipulated. In contrast, the real class includes samples from the $R_{v}F_{a}$ and $R_{v}R_{a}$ categories, whose visual content is not manipulated. As shown in Table III, this setting results in the VN training set containing 17,809 video samples per class, of which the fake class contains 9,411 samples from $F_{v}F_{a}$ and 8,398 samples from $F_{v}R_{a}$, while the real class contains 430 samples from $R_{v}R_{a}$, 430 samples from $R_{v}F_{a}$, and 16,949 samples from the external VoxCeleb1 dataset [45]. External data is used to balance the training data across classes.
Audio-only Network (AN): AN learns to detect acoustic manipulations in videos. Therefore, the training set for the AN classifier is organized by treating categories with manipulated acoustic streams as fake and categories with unmanipulated acoustic streams as real. As shown in Table II, the fake class contains all samples of $R_{v}F_{a}$ and $F_{v}F_{a}$, while the real class contains samples from $F_{v}R_{a}$ and $R_{v}R_{a}$. As shown in Table III, the training set for AN contains 9,841 samples in the fake class and 8,828 samples in the real class. For the fake class, 430 samples belong to $R_{v}F_{a}$ and 9,411 samples belong to $F_{v}F_{a}$, while for the real class, 8,398 samples belong to $F_{v}R_{a}$ and 430 samples belong to $R_{v}R_{a}$.
Audio-Visual Network (AVN): As shown in Table II, for the AVN classifier, we specify $F_{v}F_{a}$, $F_{v}R_{a}$ and $R_{v}F_{a}$ categories, where acoustic and/or visual streams are manipulated, to the fake class. However, the real class only contains video samples from the $R_{v}R_{a}$ category without acoustic and visual manipulations. As shown in Table III, the AVN training set contains 36,411 samples: 18,239 samples for the fake class and 18,172 for the real class. For the fake class, 9,411 samples belong to $F_{v}F_{a}$, 8,398 samples belong to $F_{v}R_{a}$, and 430 samples belong to $R_{v}F_{a}$. For the real class, in addition to the 430 samples from $R_{v}R_{a}$, we borrow 17,742 samples from the VoxCeleb1 dataset to balance the training data.
TABLE III: Number of training samples for different networks.
Classifier Class Samples Total Training Samples
VN Fake 17,809 35,618
Real 17,809
AN Fake 9,841 18,669
Real 8,828
AVN Fake 18,239 36,411
Real 18,172
IV-A3 Data Analysis and Preprocessing
Each classifier involves different data processing steps to ensure effective training and evaluation. The video-only network takes as input a short clip of video; therefore, we split the video into multiple short clips and extract a sequence of consecutive frames from each clip to feed into the network. To prepare the data for the audio-only network, we extract the audio track with a sampling rate of 16kHz from each video. Furthermore, we extract mel-spectrograms from these audios, which are then used as input to the audio-only network. These mel-spectral features are also used as one of the inputs to the audio-visual network. For the AV-HuBERT-based AVN, we extract the lip image frames from each video and pair them with the mel-spectral features as input to the network.
IV-B Hyperparameters
We use the Adam optimizer to train each classifier. AN (based on AST) is trained with a learning rate of 0.00001, VN (based on ViViT) is trained with a learning rate of 0.0001 and AVN (with AV-HuBERT) is trained with a learning rate of 0.002. Similarly, AVN (with AST&ViViT) is trained with a learning rate of 0.0001. The linear layer in AVTENet for combining AST-based AN, ViViT-based VN, and AV-HuBERT-based AVN is trained with a learning rate of 0.002, while the linear layer in AVTENet for combining AST-based AN, ViViT-based VN and AST&ViViT-based AVN is trained with a learning rate of 0.002.
IV-C Evaluation Metrics
We use precision, recall, F1-score, and accuracy to evaluate the performance of individual uni-modal/multi-modal classifiers and our proposed multi-modal AVTENet classifier. For all metrics, a higher value indicates better performance. For a given test input, a binary classifier produces a binary prediction output, either 1 (for positive) or 0 (for negative). Compared to the ground-truth label, the prediction can be true positive (TP: correct positive prediction), false positive (FP: incorrect positive prediction), true negative (TN: correct negative prediction), and false negative (FN: incorrect negative prediction).
Accuracy is the primary metric. It calculates the proportion of correct predictions out of total predictions:
$\mathrm{Accuracy}={\frac{TP+TN}{TP+TN+FP+FN}}.$ (6)
The precision score determines the capability of the model to correctly identify positive samples. It is a measure of the proportion of true positive predictions to the total number of positive predictions:
$\mathrm{Precision}={\frac{TP}{TP+FP}}.$ (7)
The recall score, also known as sensitivity or true positive rate, measures how much a model captures positive instances. It is an estimate of the proportion of true positive predictions to the actual positive instances in the dataset:
$\mathrm{Recall}={\frac{TP}{TP+FN}}.$ (8)
The F1-score is the harmonic mean of precision and recall, which combines them in one metric. It is calculated by taking into account false positives and false negatives:
[…]
IV-D Experimental Results
This section provides detailed experimental results and their analysis. Testset-I and Test-II are the main evaluation test sets of the FakeAVCeleb dataset.
IV-D1 Video-only Detection Results
TABLE IV: Detection Results of the Video-only Network.
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 1.00 0.87 0.93 0.94
Fake 0.89 1.00 0.94
faceswap-wav2lip Real 1.00 0.87 0.93 0.94
Fake 0.89 1.00 0.94
fsgan Real 0.95 0.87 0.91 0.91
Fake 0.88 0.96 0.92
fsgan-wav2lip Real 1.00 0.87 0.93 0.94
Fake 0.89 1.00 0.94
RTVC Real 0.50 0.87 0.64 0.50
Fake 0.50 0.13 0.20
wav2lip Real 0.97 0.87 0.92 0.92
Fake 0.88 0.97 0.93
Testset-I Real 0.87 0.87 0.87 0.87
Fake 0.87 0.87 0.87
Testset-II Real 0.76 0.87 0.81 0.80
Fake 0.85 0.73 0.78
The detailed results of the VN classifier on eight different test sets are shown in Table IV. We observe that the VN classifier has better performance on the test sets containing visual manipulations because the VN classifier utilizes visual features to detect forgeries in videos. The accuracy on the faceswap, faceswap-wav2lip, fsgan, fsgan-wav2lip, and wav2lip test sets is all above 0.9, except for RTVC, which achieves an accuracy of 0.5. Table I evidently shows that, except for RTVC, all these test sets contain visual manipulations, making VN perform better on these test sets. The poor performance on RTVC is due to the lack of visual manipulations. Here, it is important to highlight that besides visual manipulations, both faceswap-wav2lip and fsgan-wav2lip also contain acoustic manipulation, but the VN classifier only detects visual manipulation in these instances. Furthermore, the accuracy of the two main test sets (0.87 for Test set-I and 0.80 for Testset-II) is relatively low compared to other test sets. This is because, for both test sets, the fake class contains video instances with or without video manipulation. The conclusion drawn from this experiment is that the VN classifier performs well when the video is visually manipulated but fails completely when the video only contains acoustic manipulation.
TABLE V: Detection Results of the Audio-only Network.
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 0.50 1.00 0.67 0.50
Fake 0.00 0.00 0.00
faceswap-wav2lip Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
fsgan Real 0.50 1.00 0.67 0.50
Fake 0.00 0.00 0.00
fsgan-wav2lip Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
RTVC Real 0.97 1.00 0.99 0.99
Fake 1.00 0.97 0.99
wav2lip Real 0.69 1.00 0.82 0.78
Fake 1.00 0.56 0.72
Testset-I Real 0.71 1.00 0.83 0.80
Fake 1.00 0.60 0.75
Testset-II Real 0.76 1.00 0.86 0.84
Fake 1.00 0.69 0.81
IV-D2 Audio-only Detection Results
The AN classifier is a uni-modal method that uses acoustic features to detect whether a video is real or fake. The detailed results of AN on eight different test sets are shown in Table V. We observe a similar pattern to the results of the VN classifier in Table IV. The detection accuracy is almost perfect on the faceswap-wav2lip, fsgan-wav2lip, and RTVC test sets (all containing acoustic manipulation in the fake class). But AN fails completely on the faceswap and fsgan test sets, where fake instances are only visually manipulated while the audio tracks are all real. Furthermore, the detection accuracies of Testset-I and Testset-II (0.80 and 0.84, respectively) are lower than the strong performance of the other test sets. This is because, for both test sets, the fake class contains video instances with or without acoustic manipulation. The results of this experiment show that the AN classifier can detect acoustically manipulated videos well, but cannot detect fake videos containing only visual manipulations.
TABLE VI: Detection Results of the AV-HuBERT-based Audio-Visual Network.
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 0.79 0.96 0.86 0.85
Fake 0.95 0.74 0.83
faceswap-wav2lip Real 1.00 0.96 0.98 0.98
Fake 0.96 1.00 0.98
fsgan Real 0.91 0.96 0.93 0.93
Fake 0.95 0.90 0.93
fsgan-wav2lip Real 1.00 0.96 0.98 0.98
Fake 0.96 1.00 0.98
RTVC Real 0.96 0.96 0.96 0.96
Fake 0.96 0.96 0.96
wav2lip Real 1.00 0.96 0.98 0.98
Fake 0.96 1.00 0.98
Testset-I Real 0.92 0.96 0.94 0.94
Fake 0.96 0.91 0.93
Testset-II Real 1.00 0.96 0.98 0.98
Fake 0.96 1.00 0.98
IV-D3 Audio-Visual Detection Results
Compared to the VN and AN classifiers, which can only detect visual and acoustic manipulations in fake videos, respectively, the AVN classifier utilizes both acoustic and visual information to detect forgeries in videos. Table VI shows the detailed results of the AV-HuBERT-based AVN classifier on eight different test sets. Several observations can be drawn from the table. First, the AVN classifier achieves good detection performance on all test sets except the faceswap test set. One possible reason is that the faceswap test set only involves visual manipulation, and the AV-HuBERT feature extractor only extracts visual features from lip images, thereby losing information outside the lip region. Second, although the AV-HuBERT-based AVN classifier performs well on the test sets containing manipulation in a single modality, the performance may be slightly lower than audio-only and video-only classifiers focusing on one modality. For example, the AVN classifier achieves an accuracy of 0.96 on RTVC (containing only acoustic manipulation), but the AN classifier achieves an accuracy of 0.99 on the same test (see Table V). Third, the AV-HuBERT-based AVN classifier achieves accuracy of 0.94 and 0.98 on Testset-I and Testset-II, whose fake instances contain various manipulations in acoustic and visual streams, outperforming the VN classifier (cf. 0.87 and 0.80 in Table IV) and the AN classifier (cf. 0.80 and 0.84 in Table V). Therefore, the insight gleaned from this experiment is that the AVN classifier successfully detects forgeries in videos by taking into account both acoustic and visual information, generally outperforming VN and AN classifiers that focus on uni-modal visual or acoustic manipulation.
TABLE VII: Detection Results of $AVTENet_{mv}$ (Majority Voting).
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 0.80 1.00 0.89 0.87
Fake 1.00 0.74 0.85
faceswap-wav2lip Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
fsgan Real 0.88 1.00 0.93 0.93
Fake 1.00 0.86 0.92
fsgan-wav2lip Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
RTVC Real 0.96 1.00 0.98 0.98
Fake 1.00 0.96 0.98
wav2lip Real 0.99 1.00 0.99 0.99
Fake 1.00 0.99 0.99
Testset-I Real 0.92 1.00 0.96 0.96
Fake 1.00 0.91 0.96
Testset-II Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
TABLE VIII: Detection Results of $AVTENet_{asf}$ (Average Score fusion).
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 0.79 1.00 0.88 0.86
Fake 1.00 0.73 0.84
faceswap-wav2lip Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
fsgan Real 0.88 1.00 0.93 0.93
Fake 1.00 0.86 0.92
fsgan-wav2lip Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
RTVC Real 0.96 1.00 0.98 0.98
Fake 1.00 0.96 0.98
wav2lip Real 1.00 0.99 0.99 0.99
Fake 0.99 1.00 0.99
Testset-I Real 0.92 1.00 0.96 0.96
Fake 1.00 0.91 0.96
Testset-II Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
IV-D4 Detection Results of AVTENet
AVTENet is an ensemble of the above VN, AN, and AVN networks. We compare four fusion strategies, namely majority voting (mv), average score fusion (asf), score fusion (sf), and feature fusion (ff). The resulting models are denoted as $AVTENet_{mv}$, $AVTENet_{asf}$, $AVTENet_{sf}$, and $AVTENet_{ff}$, respectively. AVN is based on AV-HuBERT.
$\mathbf{AVTENet_{mv}}$: Table VII shows the detailed results of $AVTENet_{mv}$ on eight different test sets. $AVTENet_{mv}$ achieves perfect accuracy of 1.00 on faceswap-wav2lip, fsgan-wav2lip, and Testset-II and high accuracy of 0.99, 0.98, and 0.96 on wav2lip, RTVC, and Testset-I. This result confirms the advantages of integrating VN, AN, and AVN networks in fully considering visual and acoustic information in fake video detection. However, $AVTENet_{mv}$ achieves relatively low performances on faceswap and fsgan. As explained before, this may be because the faceswap and fsgan test sets only involve visual manipulations, while the AV-HuBERT feature extractor ignores visual information outside the lip region. Comparing Table VII with Tables IV, V, and VI, $AVTENet_{mv}$ outperforms all its component networks in all test conditions except for VN on faceswap and AN on RTVC. This result shows that the ensemble $AVTENet_{mv}$ comes with some tradeoffs in fake detection compared to modality-specific classifiers operating on the corresponding single-modality manipulation. Finally, it is worth mentioning again that $AVTENet_{mv}$ outperforms all its component networks on the two main tests, Testset-I and Testset-II.
$\mathbf{AVTENet_{asf}}$: Table VIII shows the detailed results of $AVTENet_{asf}$ on eight different test sets. Comparing Table VIII and Table VII, we can see that $AVTENet_{asf}$ performs almost the same as $AVTENet_{mv}$, except that the former achieves a slightly lower accuracy than the latter on the faceswap test set (0.86 vs 0.87).
TABLE IX: Detection results of $AVTENet_{sf}$ (Score fusion).
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 0.93 0.91 0.92 0.92
Fake 0.92 0.93 0.92
faceswap-wav2lip Real 1.00 0.91 0.96 0.96
Fake 0.92 1.00 0.96
fsgan Real 0.93 0.91 0.92 0.92
Fake 0.92 0.93 0.92
fsgan-wav2lip Real 1.00 0.91 0.96 0.96
Fake 0.92 1.00 0.96
RTVC Real 0.96 0.91 0.93 0.94
Fake 0.92 0.96 0.94
wav2lip Real 1.00 0.91 0.96 0.96
Fake 0.92 1.00 0.96
Testset-I Real 0.96 0.91 0.93 0.94
Fake 0.92 0.96 0.94
Testset-II Real 1.00 0.91 0.96 0.96
Fake 0.92 1.00 0.96
TABLE X: Detection Results of $AVTENet_{ff}$ (Feature fusion).
Test-Set type Class Precision Recall F1-Score Accuracy
faceswap Real 0.93 0.97 0.95 0.95
Fake 0.97 0.93 0.95
faceswap-wav2lip Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
fsgan Real 0.93 0.97 0.95 0.95
Fake 0.97 0.93 0.95
fsgan-wav2lip Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
RTVC Real 0.96 0.97 0.96 0.96
Fake 0.97 0.96 0.96
wav2lip Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
Testset-I Real 0.96 0.97 0.96 0.96
Fake 0.97 0.96 0.96
Testset-II Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
$\mathbf{AVTENet_{sf}}$: Table IX shows the detailed results of $AVTENet_{sf}$ on eight different test sets. Comparing Table IX with Tables VII and VIII, we can see that $AVTENet_{sf}$ performs more stably than $AVTENet_{mv}$ and $AVTENet_{asf}$ across 8 test sets. The stability may be attributed to the additional linear layer taking as input the scores output by the VN, AN, and AVN networks, as it learns to balance the contributions of these three component networks. However, although $AVTENet_{sf}$ outperforms $AVTENet_{mv}$ and $AVTENet_{asf}$ on the faceswap test set (0.92 vs 0.87 and 0.86), it performs worse than $AVTENet_{mv}$ and $AVTENet_{asf}$ on all the other test sets.
$\mathbf{AVTENet_{ff}}$: Table X shows the detailed results of $AVTENet_{ff}$ on eight different test sets. We can see that $AVTENet_{ff}$ performs best among four ensemble models on faceswap, fsgan, and Testset-II, and achieves near-perfect performance on faceswap-wav2lip, fsgan-wav2lip, wav2lip, and Testset-II. While the other ensemble models may excel in performance on specific test sets, $AVTENet_{ff}$ maintains stable and reliable performance on all eight test sets. The good performance may be attributed to the additional linear layer that takes as input the embeddings of the penultimate layers of the VN, AN, and AVN networks, as it learns to balance the contributions of these three component networks. In summary, feature fusion is the best fusion strategy in this work.
Figure 6: Performance comparison of our models on two main test sets, Testset-I and Testset-II.
Figure 7: The comparison of our proposed approach ($AVTENet_{ff}$) with other existing unimodal, multimodal, fusion, and ensemble benchmarks.
Comparison of Our Models
Fig. 6 depicts the accuracy of the above seven models evaluated on two main test sets, Tesetset-I and Testset-II. The video-only model (VN) is good at identifying visual manipulations in videos. Likewise, the audio-only model (AN) is good at identifying manipulated acoustic content in videos. But they only focus on a single modality. In contrast to VN and AN, the audio-visual model (AVN) and all variants of AVTENet are able to identify acoustic and visual manipulations in videos. Although $AVTENet_{mv}$ and $AVTENet_{asf}$ have higher accuracy than $AVTENet_{ff}$ on Testset-II, the latter shows higher stability in detecting various manipulation types in videos.
TABLE XI: Performance comparison of AVTENet (AST_ViViT_AST&ViViT) and AVTENet (AST_ViViT_AV-HuBERT).
Models
Ensemble
Strategies
faceswap
faceswap-
wav2lip
fsgan
fsgan-
wav2lip
RTVC wav2lip Testset-I Testset-II
AVTENet (AST_ViViT_AST&ViViT) $AVTENet_{mv}$ 0.96 0.99 0.94 0.99 0.61 0.98 0.87 0.94
$AVTENet_{asf}$ 0.96 0.99 0.93 0.99 0.63 0.98 0.88 0.94
$AVTENet_{sf}$ 0.91 0.93 0.88 0.93 0.52 0.91 0.80 0.87
$AVTENet_{ff}$ 0.92 0.97 0.88 0.96 0.52 0.90 0.80 0.87
AVTENet (AST_ViViT_AV-HuBERT) $AVTENet_{mv}$ 0.87 1.00 0.93 1.00 0.98 0.99 0.96 1.00
$AVTENet_{asf}$ 0.86 1.00 0.93 1.00 0.98 0.99 0.96 1.00
$AVTENet_{sf}$ 0.92 0.96 0.92 0.96 0.94 0.96 0.94 0.96
$AVTENet_{ff}$ 0.95 0.99 0.95 0.99 0.96 0.99 0.96 0.99
TABLE XII: Performance comparison of our Ensemble model and several existing uni-modal, multi-modal, fusion, and ensemble models. DFD in the first column refers to the deepfake detection method. “V”, “A” and “AV” stand for visual, audio and audio-visual modalities, respectively.
DFD Method Model Modality Class Precision Recall F1-score Accuracy
Unimodal[56] VGG16 V Real 0.6935 0.8966 0.7821 0.8103
Fake 0.8724 0.7750 0.8208
Unimodal[56] Xception A Real 0.8750 0.6087 0.7179 0.7626
Fake 0.7033 0.9143 0.7950
Unimodal [57] LipForensics V Real 0.70 0.91 0.80 0.76
Fake 0.88 0.61 0.72
Ensemble (Soft-Voting)[56] VGG16 AV Real 0.6935 0.8966 0.7821 0.7804
Fake 0.8948 0.6894 0.7788
Ensemble (Hard-Voting)[56] VGG16 AV Real 0.6935 0.8966 0.7821 0.7804
Fake 0.8948 0.6894 0.7788
Multimodal-1[56] Multimodal-1 AV Real 0.000 0.000 0.000 0.5000
Fake 0.496 1.000 0.663
Multimodal-2[56] Multimodal-2 AV Real 0.710 0.587 0.643 0.674
Fake 0.648 0.760 0.700
Multimodal-3[56] CDCN AV Real 0.500 0.068 0.120 0.515
Fake 0.500 0.940 0.651
Multimodal-4[16] Not-made-for-each-other AV Real 0.62 0.99 0.76 0.69
Fake 0.94 0.40 0.57
Multimodal [10] Ensemble AV Real 0.83 0.99 0.90 0.89
Fake 0.98 0.80 0.88
Multimodal [11] AV-Lip-Sync AV Real 0.93 0.96 0.94 0.94
Fake 0.96 0.93 0.94
Modality Mixing [58] Multimodaltrace AV Real - - - 0.929
Fake - - -
Ensemble[59] AVFakeNet AV Real - - - 0.934
Fake - - -
Fusion[60] AVoiD-DF AV Real - - - 0.837
Fake - - -
Fusion[61] MIS-AVioDD AV Real - - - 0.962
Fake - - -
Ensemble (ours) $\mathbf{AVTENet_{ff}}$ AV Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
IV-D5 Comparison of AVTENet with two different AVNs
Our proposed AVTENet model is an ensemble of an AST-based AN model, a ViViT-based VN model, and an AV-HuBERT-based AVN model. Although $AVTENet_{ff}$ performs well on all test sets, its performances on the faceswap and fsgan test sets are relatively low, where the fake instances are only visually manipulated while audio tracks are real. We speculate that the reason may be that the AV-HuBERT feature extractor in AVN only extracts visual information from lip images and ignores information outside the lip region. Therefore, we also implement another AVN based on AST and ViViT. Here, we compare AVTENet (AST_ViViT_AV-HuBERT) and AVTENet (AST_ViViT_AST&ViViT). From Table XI, it is clear that although the AST_ViViT_AST&ViViT-based $AVTENet_{mv}$ and $AVTENet_{asf}$ models do slightly improve the performance on the faceswap test set, the AST_ViViT_AST&ViViT-based AVTENet models perform worse than the AST_ViViT_AV-HuBERT-based AVTENet models. The reason may be that since the AST and ViViT models are already used in the AN and VN models, they cannot provide additional information in the AVN model. Surprisingly, the AST_ViViT_AST&ViViT-based AVTENet models almost fail on the RTVC test set. To fully utilize face images in the AVN model, we may need to apply new pre-trained visual models.
IV-D6 Comparison of AVTENet and other Models
Finally, as shown in Fig. 7 and Table XII, we conduct a thorough comparison of our proposed $AVTENet_{ff}$ model with various existing uni-modal, multi-modal, fusion, and ensemble methods. Several uni-modal, multi-modal and ensemble methods were presented in [56], but most of them failed to detect acoustic and visual manipulations in videos well. VGG16, a uni-modal approach exclusively trained on visual data, outperforms all other uni-modal approaches reported in [56] for video forgery detection with an accuracy of 0.81. Xception, another uni-modal method reported in [56], is trained exclusively on the audio modality and achieves an accuracy of only 0.76. Similarly, Lip Forensics [35] exploits visual lip movements to detect forgeries in videos with an accuracy of 0.76. Although Ensemble (soft-voting and hard-voting), Multi-modal-1, Multi-modal-2, and Multi-modal-3 utilize both audio and visual modalities for deepfake detection, their performance is still unsatisfactory. Not-made-for-each other [16] is an audio-visual dissonance-based approach that detects video forgeries with an accuracy of 0.69. The multi-modal ensemble approach proposed in [10] utilizes three CNN-based classifiers to identify forgeries and achieves an accuracy of 0.89. Another audio-visual approach proposed in [11] successfully identifies fake videos using lip synchronization, and the model achieves an accuracy of 0.94. Moreover, Multimodaltrace is an ensemble approach proposed in [58], which achieves an accuracy of 0.929. AVFakeNet [59] is also an ensemble approach using audio and video transformer encoders, which achieves an accuracy of 0.934. AVoiD-DF [60] is another audio-visual fusion approach, but it only achieves an accuracy of 0.837. MIS-AVioDD [61] is also a fusion approach that jointly uses the modality invariant and modality-specific representations to detect the audio-visual forgery and achieves an accuracy of 0.962. From Table XII, it is evident that our ensemble approach outperforms other existing models, achieving an accuracy of 0.99 on Testset-II of the FakeAVCeleb dataset, which is a new SOTA performance.
V Conclusions and Future Work
In this work, we demonstrated the effectiveness of ensemble learning and studied deepfake video detection from an audio-visual perspective. We proposed a novel audio-visual transformer-based ensemble learning solution for effective and scalable video forgery detection, especially for videos containing acoustic and visual manipulations, which cannot rely on existing methods because they can only detect single-modal manipulations or lack training diversity. Specifically, we have devised a transformer-based model with a powerful detection ability due to its sequence modeling, parallelization, and attention mechanisms. We have illustrated the performances of our multi-modal transformer-based ensemble model on several manipulation techniques and compared it with various existing uni-modal, multi-modal, fusion, and ensemble models. Our model achieves state-of-the-art performance on the FakeAVCeleb dataset. It should be emphasized that the proposed solution is extensible via self-supervised learning models with large amounts of training data. To further improve upon the results presented in the paper, we plan to advance this research by integrating more recent self-supervised learning models.